OpenStudy (anonymous):

1. Which binomial is a factor of f(x) = x^4 + 7x^3 - 7x^2 - 43x + 42? a. (x - 7) b. (x - 3) c. (x - 1) d. (x + 2) 2. Which is the factored form of x3 + 6x2 25x + 18 = 0? a. (x + 1)(x - 3)(x - 6) = 0 b.(x + 1)(x 2)(x + 9) = 0 c. (x - 1)(x - 3)(x - 6) = 0 d. (x - 1)(x - 2)(x + 9) = 0

jhonyy9 (jhonyy9):

x*4+7x*2(x-1)-42x-x+42=x*4+7x*2(x-1)-42(x-1)+x=x*4+x+(x-1)(7x*2 -42)= =x(x*3 +1)+(x-1)(7x*2 -42)=x(x-1)(x*2+x-1)+(x-1)(7x*2-42)= =(x-1)(x(x*2+x-1)+7(x*2-6)) --- resulted c.

jhonyy9 (jhonyy9):

2. please wtite the missed signes in equation after - next - 6x*2 and in point b. (x+1)(x ??? 2)(x+9)